QUESTION 7

  1. Shares of a closed-end fund are trading at a 6% premium to its $10 net asset value. The current market price of the fund's shares are closest to:

    a. $9.40.

    b. $10.06.

    c. $10.60.

    d. $16.00.

QUESTION 8

  1. Pacific Fund invests exclusively in Treasury bills and other debt securities with very short-term maturities. Pacific can most likely be characterized as a:

    a. Growth fund.

    b. Fixed-income fund.

    c. Money market fund.

    d. Balanced fund.
Answer #1

7

Price = NAV*(1+premium%) = 10*(1+0.06) = 10.6

phpVpPS6a.png